OpenStudy (simplymarie_x):

A peer is having difficulty finding angle measures using trigonometric ratios in a right triangle. Create an example to demonstrate how to do this and be sure to address using the reciprocal relationships (cosecant, secant, cotangent) too.

OpenStudy (baru):

i think the question is just asking you to take any right triangle, mark the sides, take a reference angles, indentify the hypotenuse, and write all the trignometric ratios for that triangle ex: \(sin(\theta\))= AB/AC

OpenStudy (simplymarie_x):

Possibly but I don't understand cosecant and all that jazz

OpenStudy (simplymarie_x):

I think I'll just forfeit this question, and write down just the formulas. A couple points off won't hurt me lol

OpenStudy (baru):

hey you already know sin and cos dont you, cosec = 1/sin ( its just the reciprocal of sin) sec = 1/cos

OpenStudy (simplymarie_x):

Oh you're right. Thanks for the help though :-)

OpenStudy (baru):

cosec = AC/AB like literally just flip sin upside down, as simple as that.

OpenStudy (baru):

you're welcome :)